OpenStudy (yinshy):

What is the exact area of a circular patio with a radius of 17 ft? A. 8.5π ft² B. 17π ft² C. 72.25π ft² D. 289π ft²

OpenStudy (calculusxy):

Area = \(\pi\)\(r^2\) This is in terms of pi. So just square the radius and include it with the pi

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (calculusxy):

Area = \(\pi\)\(r^2\) Area = \(\pi\)\((17^2)\)

OpenStudy (yinshy):

?

OpenStudy (calculusxy):

What is \(17^2\)?

OpenStudy (yinshy):

34

OpenStudy (calculusxy):

I didn't say 17 x 2. I said to evaluate \(17^2\) which is the same thing as \(17 \times 17\).

OpenStudy (anonymous):

YOU CAN DO IT

OpenStudy (yinshy):

oh oops lol

OpenStudy (yinshy):

289

OpenStudy (calculusxy):

Good. Now attach the pi symbol with it.

OpenStudy (yinshy):

D

OpenStudy (calculusxy):

Correct!